The sum of two numbers is 47, and their difference is 15. What is the largest number?
Answer by truly_smart_one(1)

x+y=47 x-y=15 ==>> so, add those two together, to make 2x=62 then divide 62 by 2 which is x=31. Then the equation is 31+y=47 47-31=16 so y is 16 and x is 31. 31 is the largest answer